OpenStudy (anonymous):

6x^2 − 19x + 3 = 0 is 1/6. What is the other solution?

OpenStudy (anonymous):

\[6x^2 - 19x + 3 = 0\] is \[\frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

lol

OpenStudy (zzr0ck3r):

6*3 = 18 and -18+ -1=-19 so 6x^2-19x+3 = 6x^2-18x-x+3 = 6x(x-3)-1(x-3) = (6x-1)(x-3) = 0 so 6x-1=0 implies x = 1/6 and x-3 = 0 implies x = ?

OpenStudy (zzr0ck3r):

sorry its late:)

OpenStudy (anonymous):

I was a bit confused when you took it off xD

OpenStudy (anonymous):

Ok so let me review that one sec...

OpenStudy (anonymous):

one zero is \(\frac{1}{6}\) so the factor is \(6x-1\)

OpenStudy (anonymous):

OH so you basically just made 2 terms out of -19x right?

OpenStudy (anonymous):

put \[6x^2-19x+3=(6x-1)(something)=0\]

OpenStudy (anonymous):

x = 3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OHHH ok I see it now, after reviewing the..er...scatter zz made :)

OpenStudy (zzr0ck3r):

the method I used is sometimes called the AC method. you should Google that and practice it. imho

OpenStudy (anonymous):

haha no problem I will :)

OpenStudy (anonymous):

The thirst for knowledge!